OpenStudy (anonymous):

I know the answer, just teach me how to work it out. x2 = 14x – 49

OpenStudy (karatechopper):

ok well first since this is a quadratic equation, you wanna make the equation equal 0

OpenStudy (karatechopper):

So you would do this step first... x2-x2=-x2+14x-49

OpenStudy (anonymous):

Key pattern for this one: a^2-2ab+b^2=(a-b)^2 Once we recall this, it's easy. \[x^2 = 14x – 49 \implies x^2 -14x + 49=0=(x-7)^2 \implies x=7\]

OpenStudy (karatechopper):

making that become, 0=-x2+14x-49 only there is one problem..a number with an exponent can never be negative! So we change the signs of everything! like this! 0=x2-14x+49 NOTE- i did not put a negative in front of zero because it would make no different..

OpenStudy (karatechopper):

Now, do you know how to factor things? @destinyshaw

OpenStudy (anonymous):

Yes I do

OpenStudy (karatechopper):

ok then go ahead and factor it out for me!

OpenStudy (anonymous):

@AnimalAin just did it for me... but I just looked at my work, and tell me what factorin gis exactly just in case it is a test question in the future